# Suppression of Fast Flavor Conversion by Red Turbulence in Supernovae

### v1

@@ Abstract
Fast flavor conversions (FFCs) in supernovae, driven by neutrino-neutrino refraction, can catastrophically equilibrate flavors and potentially affect the neutrino-driven explosion. We present a pivotal insight: matter density fluctuations characterized by red spectra ($\nu<0$), naturally arising in stratified supernova environments, can suppress such instabilities by inducing accelerated decoherence. By deriving exact analytical solutions for two-flavor evolution in red turbulent matter—where correlations grow as $t^{|\nu|}$—we uncover a novel acceleration of coherence loss. This dynamical decoherence mechanism raises an effective energy barrier against the collective growth of flavor instabilities. Translating our master-equation results into an effective damping rate for FFC linear analysis, we find that realistic red turbulence ($\nu\sim-1$, fluctuation strength $\xi_{\nu}\sim 0.1$) can elevate the FFC threshold by a factor of $\sim 3-5$, potentially stabilizing regions that would otherwise undergo explosion-killing flavor equilibration (or vice versa). Our work provides the first analytical criterion for FFC suppression in turbulent media and identifies red turbulence as a critical, physics-grounded ingredient missing from current supernova models.
@@
In this Letter, we present exact analytical solutions for neutrino evolution in red turbulent matter. Solving the non-Markovian master equation with kernel $K(t)=\xi_{\nu}(t/\tau_{c})^{|\nu|}e^{-t/\tau_{c}}$ ($\nu<0$), we discover that red turbulence generates fractional integral operators of order $1+|\nu|$, contrasting with the fractional derivatives of Kolmogorov turbulence. This leads to a key physical effect: accelerated decoherence—slower damping of flavor coherence compared to white noise of equivalent fluctuation strength.
We translate this microphysical insight to FFCs: the derived decoherence rate $\Gamma_{\text{dec}}$ acts as an effective damping term in the linear stability analysis. For realistic supernova parameters ($\nu\sim-1$, $\xi_{\nu}\sim 0.1$, $\tau_{c}\sim 0.1$ km), $\Gamma_{\text{dec}}\sim\mathcal{O}(10^{3})\,\text{km}^{-1}$ can raise the FFC threshold by a factor of 3–5, potentially stabilizing regions vulnerable to explosion-killing flavor equilibration or vice versa. Our work provides both a novel analytical framework for neutrinos in correlated media and a physics-based mechanism to resolve the tension between predicted FFC instabilities and the requirement of sustained neutrino heating for successful explosions.
Analytical Two-Flavor Solutions for Red Turbulence.—We present exact analytical solutions for neutrino evolution in turbulent matter with red spectra ($\nu<0$), where density correlations grow algebraically as $t^{|\nu|}$. The complete mathematical derivation is provided in the companion paper [1].
All quantities are expressed in natural units ($c=\hbar=1$). For red turbulence with spectral exponent $\nu<0$, the correlation function exhibits the distinctive form $K(t)\propto t^{-\nu}e^{-t/\tau_{c}}=t^{|\nu|}e^{-t/\tau_{c}}$, contrasting with the decaying correlations ($t^{-5/3}$) of Kolmogorov turbulence. The fundamental parameters are:
@@
$\xi_{\nu}$ Red turbulence strength parameter dimensionless
@@
The non-Markovian master equation for neutrino evolution in red turbulent matter is:
@@
$-i[\hat{H}_{0},\langle\hat{\rho}(t)\rangle]-g_{\nu}\int_{0}^{t}ds\,(t-s)^{|\nu|}e^{-(t-s)/\tau_{c}}[\sigma_{z},[\sigma_{z},\langle\hat{\rho}(s)\rangle]],$ (1)
where $\hat{H}_{0}=\frac{\Delta_{m}}{2}(-\cos 2\theta_{m}\sigma_{z}+\sin 2\theta_{m}\sigma_{x})$ is the effective Hamiltonian in the matter basis, with $\Delta_{m}=\sqrt{(\frac{\Delta m^{2}}{2E}\cos 2\theta-\sqrt{2}G_{F}n_{e})^{2}+(\frac{\Delta m^{2}}{2E}\sin 2\theta)^{2}}$ being the matter oscillation frequency. The kernel $K(t)=g_{\nu}t^{|\nu|}e^{-t/\tau_{c}}$ encodes the algebraic growth characteristic of red turbulence, with $g_{\nu}$ quantifying the turbulence strength [1].
@@
where $\Gamma(1-\nu)>0$ for $\nu<0$. The key mathematical distinction emerges: while Kolmogorov turbulence ($\nu=5/3$) generates fractional derivatives of order $2/3$, red turbulence ($\nu<0$) produces fractional integrals of order $1+|\nu|$. This fundamentally alters the decoherence dynamics.
@@
The perturbative approach naturally extends to three flavors via Gell-Mann matrix generalizations [9].
Asymptotic Analysis for Red Turbulence.—Red turbulence produces accelerated decoherence with distinctive scaling laws. We define dimensionless parameters: turbulence strength $\xi_{\nu}\approx(\delta n_{e}/\bar{n}_{e})^{2}$, memory parameter $\alpha=\Delta_{m}\tau_{c}$, and scaled time $x=\Delta_{m}t$. The survival probability exhibits universal scaling:
$P_{ee}(x)=\frac{1}{2}+\frac{1}{2}\cos 2\theta_{m}\,F_{\nu}(x,\alpha,\xi_{\nu})\cos(x),$ (4)
where $F_{\nu}$ is the red turbulence scaling function.
Regime I: Early-Time Accelerated Decoherence ($x\ll\alpha$): For $t\ll\tau_{c}$, red turbulence produces accelerated decoherence scaling as $x^{2+|\nu|}$:
$P_{ee}^{(I)}(x)=1-\frac{\sin^{2}2\theta_{m}}{4}x^{2}-\xi_{\nu}\,\mathcal{F}_{\text{red}}(x)+\mathcal{O}(x^{4+2|\nu|}),$ (5)
where:
$\mathcal{F}_{\text{red}}(x)\approx C_{\nu}x^{2+|\nu|}+C_{2}x^{3+|\nu|}+\cdots,$ (6)
with $C_{\nu}=[\Gamma(1-\nu)/(2+|\nu|)][1/(1+|\nu|)-1/(2+|\nu|)]>0$. For $\nu=-1$, this yields $x^{3}$ scaling, significantly faster than Kolmogorov’s $x^{5/3}$.
Regime II: Crossover at Correlation Time ($x\sim\alpha$): At $t\sim\tau_{c}$:
$F_{\nu}(x,\alpha,\xi_{\nu})=\exp\left[-\xi_{\nu}\Gamma(1-\nu)\frac{\gamma(1-\nu,x/\alpha)}{\Gamma(1-\nu)}\right]\left[1+\mathcal{O}(\xi_{\nu}^{2})\right],$ (7)
where $\gamma(s,z)$ is the incomplete Gamma function. Red turbulence shows enhanced memory effects due to the integral nature of the fractional operator.
Regime III: Late-Time Markovian Limit ($x\gg\alpha$): For $t\gg\tau_{c}$:
$F_{\nu}(x,\alpha,\xi_{\nu})\approx\exp\left(-\frac{\xi_{\nu}}{\alpha}x\right)\left[1+\mathcal{O}\left(\frac{e^{-x/\alpha}}{x^{1-\nu}}\right)\right],$ (8)
recovering exponential decay with rate $\Gamma_{\text{eff}}=\xi_{\nu}/\tau_{c}$, but with algebraic corrections decaying as $x^{-(1+|\nu|)}$.
@@
Figure 1: Electron neutrino survival probability $P_{ee}(x)$ for turbulence with different spectral indices $\nu$, demonstrating the effect of temporal correlations on decoherence. White noise ($\nu=0$, blue solid line) produces the strongest decoherence, reaching the lowest asymptotic value. Red noise with $\nu=-0.5$ (orange dashed), $\nu=-1.0$ (green dash-dot), and $\nu=-2.0$ (red dotted) shows progressively weaker decoherence as $\nu$ decreases, with higher asymptotic survival probabilities. This indicates that long-range temporal correlations (red spectra) preserve flavor coherence more effectively than uncorrelated (white) fluctuations, despite having identical integrated turbulence strength $\xi=0.1$. The small-scale regularization $\epsilon=0.01$ (we introduce a physical cutoff $\epsilon\sim\tau_{\text{diss}}$ at the dissipation scale, modifying the kernel to $K(t)\propto t^{-\nu}\exp(-\epsilon/t)$. For $\nu<0$, this regularization ensures mathematical consistency while having negligible effect on physical results) ensures convergence of all memory integrals.
@@
The key distinctions of red turbulence are: (1) Accelerated early-time scaling: $t^{2+|\nu|}$ vs. Kolmogorov’s $t^{5/3}$ (2) Fractional integrals rather than derivatives (3) Stronger memory effects at crossover (4) Faster perturbative convergence due to regular kernel at $t=0$
@@
Im[$\imaginary[\omega_{\text{FFC}}]>\Gamma_{\text{dec}}(\nu,\xi_{\nu},\tau_{c}),$ (9)
@@
To quantify the potential suppression, we evaluate $\Gamma_{\text{dec}}$ using turbulence parameters from supernova simulations. Multi-dimensional studies show significant density fluctuations in the gain region, with relative amplitudes $\delta n_{e}/\bar{n}_{e}\sim 40\%$ yielding $\xi_{\nu}\sim 0.16$ [13]. The turbulence correlation scale in stratified regions like the proto-neutron star atmosphere is typically $L_{c}\sim 0.1-1$ km, giving $\tau_{c}\sim 0.1$ km in natural units [6]. For stratified turbulence, spectral exponents $\nu$ range from 1.0-1.0 (strong stratification) to 0.5-0.5 (moderate stratification) [5].
For the representative case $\nu=-1$, our solution gives $\Gamma_{\text{dec}}\sim\xi_{\nu}\tau_{c}\omega^{2}$ when $\omega\tau_{c}\ll 1$, where $\omega$ characterizes the neutrino oscillation frequency. In the gain region, the neutrino-neutrino potential $\mu_{\nu\nu}\sim 10^{2}-10^{4}$ km-1. Taking $\omega\sim\mu_{\nu\nu}\sim 10^{3}$ km-1 as a typical value, we obtain:
$\Gamma_{\text{dec}}\sim 0.16\times 0.1\,\text{km}\times(10^{3}\,\text{km}^{-1})^{2}\sim 1.6\times 10^{4}\,\text{km}^{-1}.$ (10)
This damping rate exceeds typical FFC growth rates in the gain region ($\sim 10^{1}-10^{3}$ km-1) by approximately an order of magnitude. The instability threshold is thus raised by a factor of 3-10, with the conservative estimate being a 3-5 fold enhancement.
@@
For practical implementation, we propose that future supernova simulations extract turbulence parameters ($\nu$, $\xi_{\nu}$, $\tau_{c}$) along neutrino trajectories and apply the criterion Im[$\imaginary[\omega_{\text{FFC}}]>\Gamma_{\text{dec}}$ to determine FFC viability region-by-region. This provides a physics-based approach to assess whether and where FFCs might develop in realistic turbulent environments.
@@
The key physical discovery is that red turbulence induces accelerated decoherence: for the same integrated fluctuation strength $\xi_{\nu}$, flavor coherence persists longer than in white-noise environments but decays differently than in Kolmogorov-type turbulence. This decoherence behavior emerges from the algebraically growing temporal correlations $K(t)\sim t^{|\nu|}$ characteristic of stratified astrophysical plasmas.
Applying our results to the supernova fast flavor conversion problem, we find that the derived decoherence rate $\Gamma_{\text{dec}}(\nu,\xi_{\nu},\tau_{c})$ can act as an effective damping mechanism in the linear stability analysis. For realistic parameters in supernova gain regions and proto-neutron star atmospheres ($\nu\sim-1$, $\xi_{\nu}\sim 0.1$, $\tau_{c}\sim 0.1$ km), this damping rate is $\mathcal{O}(10^{3}-10^{4})$ km-1—comparable to or exceeding typical FFC growth rates. Consequently, red turbulence can raise the FFC instability threshold by a factor of 3-5, potentially suppressing explosion-killing flavor equilibration in critical regions.

### v2

@@ Abstract
Fast flavor conversions (FFCs) in supernovae, driven by neutrino-neutrino refraction, can catastrophically equilibrate flavors and potentially affect the neutrino-driven explosion. We present a pivotal insight: matter density fluctuations characterized by extremely red spectra ($\nu<0$, where $|\nu|\equiv-\nu>0$) naturally arising in stratified supernova environments, can suppress such instabilities by inducing accelerated decoherence. By deriving exact analytical solutions for two-flavor evolution in red turbulent matter—where correlations grow as $t^{|\nu|}$—we uncover a novel acceleration of coherence loss. This dynamical decoherence mechanism raises an effective energy barrier against the collective growth of flavor instabilities. Translating our master-equation results into an effective damping rate for FFC linear analysis, we find that realistic red turbulence ($\nu\sim-1$, fluctuation strength $\xi\sim 0.1$) can elevate the FFC threshold by a factor of $\sim 3-5$, potentially stabilizing regions that would otherwise undergo explosion-killing flavor equilibration (or vice versa). Our work provides the first analytical criterion for FFC suppression in turbulent media and identifies red turbulence as a critical, physics-grounded ingredient missing from current supernova models.
@@
In this Letter, we present exact analytical solutions for neutrino evolution in red turbulent matter. Solving the non-Markovian master equation with kernel $K(t)=\xi(t/\tau_{c})^{|\nu|}e^{-t/\tau_{c}}$ ($\nu<0$), we discover that red turbulence generates fractional integral operators of order $1+|\nu|$, contrasting with the fractional derivatives of Kolmogorov turbulence. This leads to a key physical effect: accelerated decoherence—slower damping of flavor coherence compared to white noise of equivalent fluctuation strength.
We translate this microphysical insight to FFCs: the derived decoherence rate $\Gamma_{\text{dec}}$ acts as an effective damping term in the linear stability analysis. For realistic supernova parameters ($\nu\sim-1$, $\xi\sim 0.1$, $\tau_{c}\sim 0.1$ km), the turbulence can raise the FFC threshold by a factor of 3–5, potentially stabilizing regions vulnerable to explosion-killing flavor equilibration or vice versa. Our work provides both a novel analytical framework for neutrinos in correlated media and a physics-based mechanism to resolve the tension between predicted FFC instabilities and the requirement of sustained neutrino heating for successful explosions.
Analytical Two-Flavor Solutions for Red Turbulence.—We present exact analytical solutions for neutrino evolution in turbulent matter with extremely red spectra ($\nu<0$, $|\nu|\equiv-\nu>0$), where density correlations grow algebraically as $t^{|\nu|}$. This regime, characterized by negative spectral exponents, emerges naturally in stratified supernova environments and contrasts with moderately red spectra ($0<\nu<1$) studied in our companion paper [1].
All quantities are expressed in natural units ($c=\hbar=1$). For red turbulence with spectral exponent $\nu<0$, the correlation function exhibits the distinctive form. Using the Nakajima-Zwanzig projection operator method [4], the exact memory kernel for classical Gaussian noise is derived as
$\mathcal{K}(\tau)=\langle V(0)V(\tau)\rangle\,e^{-\Phi(\tau)[\sigma_{z},[\sigma_{z},\cdot]]}\,[\sigma_{z},[\sigma_{z},\cdot]],$ (1)
where $\Phi(\tau)=\frac{1}{2}\!\int_{0}^{\tau}\!d\tau_{1}\!\int_{0}^{\tau_{1}}\!d\tau_{2}\,\langle V(\tau_{1})V(\tau_{2})\rangle$, contrasting with the decaying correlations ($t^{-5/3}$) of Kolmogorov turbulence. The fundamental parameters are:
@@
$\xi$ Red turbulence strength parameter dimensionless
@@
The non-Markovian master equation for neutrino evolution in extremely red turbulent matter ($\nu<0$) is:
@@
$-i[\hat{H}_{0},\langle\hat{\rho}(t)\rangle]-4\xi\int_{0}^{t}ds\,(t-s)^{|\nu|}e^{-(t-s)/\tau_{c}}[\sigma_{z},[\sigma_{z},\langle\hat{\rho}(s)\rangle]],$ (2)
where $\hat{H}_{0}=\frac{\Delta_{m}}{2}(-\cos 2\theta_{m}\sigma_{z}+\sin 2\theta_{m}\sigma_{x})$ is the effective Hamiltonian in the matter basis, with $\Delta_{m}=\sqrt{(\frac{\Delta m^{2}}{2E}\cos 2\theta-\sqrt{2}G_{F}n_{e})^{2}+(\frac{\Delta m^{2}}{2E}\sin 2\theta)^{2}}$ being the matter oscillation frequency. The kernel $K(t)=\xi t^{|\nu|}e^{-t/\tau_{c}}$ encodes the algebraic growth characteristic of red turbulence, with $\xi$ quantifying the turbulence strength [1].
@@
where $\Gamma(1+|\nu|)>0$ for $\nu<0$. The key mathematical distinction emerges: while Kolmogorov turbulence ($\nu=5/3$) generates fractional derivatives of order $2/3$, red turbulence ($\nu<0$) produces fractional integrals of order $1+|\nu|$. This fundamentally alters the decoherence dynamics.
@@
The approach naturally extends to three flavors via Gell-Mann matrix generalizations [10].
I Universal Scaling and Exact Solution
The neutrino flavor evolution in red turbulence exhibits a remarkable universal scaling that transcends specific parameter choices. Introducing the dimensionless variables
$x$ $\equiv\Delta_{m}t,\quad\alpha\equiv\Delta_{m}\tau_{c},$ (5)
$\chi_{\nu}$ $\equiv 4\xi\tau_{c}^{|\nu|+1}\Delta_{m}^{|\nu|+2}\Gamma(1+|\nu|),$ (6)
$\xi$ $\equiv\frac{G_{F}^{2}\langle(\delta n_{e})^{2}\rangle}{2\Delta_{m}^{2}},$ (7)
where $\langle(\delta n_{e})^{2}\rangle$ is the mean-square density fluctuation. The survival probability collapses onto a universal curve
$P_{ee}(x)=\frac{1}{2}+\frac{1}{2}\cos 2\theta_{m}\,\mathcal{F}_{\nu}(x;\alpha,\chi_{\nu}),$ (8)
where the scaling function $\mathcal{F}_{\nu}$ satisfies the fractional master equation
$\mathcal{F}_{\nu}^{(|\nu|+2)}(x)-i\mathcal{F}_{\nu}^{(|\nu|+1)}(x)-\chi_{\nu}\mathcal{F}_{\nu}(x)=0,\qquad\mathcal{F}_{\nu}(0)=1.$ (9)
Here, $\mathcal{F}_{\nu}^{(n)}$ denotes the $n$th derivative, and the imaginary unit $i$ reflects the coherent oscillation phase.
I.1 Mittag-Leffler Exact Solution
Equation (9) admits an elegant closed-form solution in terms of multivariate Mittag-Leffler functions
$\mathcal{F}_{\nu}(x)=E_{(|\nu|+2,\,|\nu|+1)}\Bigl(ix^{|\nu|+2},\;-\chi_{\nu}x^{|\nu|+2}\Bigr),$ (10)
where the $(p,q)$-Mittag-Leffler function is defined by the series
$E_{(p,q)}(a,b)=\sum_{n=0}^{\infty}\frac{a^{n}x^{pn}}{(pn)!}\;{}_{1}F_{1}\!\left(\begin{array}[]{c}1\\ qn+1\end{array};bx^{q}\right).$ (11)
The confluent hypergeometric function $_{1}F_{1}$ in Eq. (11) encodes the interplay between coherent oscillations (parameter $a$) and stochastic damping (parameter $b$). This exact representation reveals how temporal correlations—encoded in the fractional derivative order $|\nu|+2$—fundamentally reshape the decoherence dynamics beyond the Markovian paradigm.
@@ I.1 Mittag-Leffler Exact Solution
Figure 1: Electron neutrino survival probability $P_{ee}(x)$ for turbulence with different spectral indices $\nu$, demonstrating the effect of temporal correlations on decoherence. White noise ($\nu=0$, blue solid line) produces the strongest decoherence, reaching the lowest asymptotic value. Red noise with $\nu=-0.5$ (orange dashed), $\nu=-1.0$ (green dash-dot), and $\nu=-2.0$ (red dotted) shows progressively weaker decoherence as $\nu$ decreases, with higher asymptotic survival probabilities. This indicates that long-range temporal correlations (red spectra) preserve flavor coherence more effectively than uncorrelated (white) fluctuations, despite having identical integrated turbulence strength $\xi=0.1$. For $\nu<0$, this regularization ensures mathematical consistency while having negligible effect on physical results) ensures convergence of all memory integrals.
@@
The key distinctions of red turbulence are: (1) Accelerated early-time scaling: $t^{2+|\nu|}$ vs. Kolmogorov’s $t^{5/3}$ (2) Fractional integrals rather than derivatives (3) Stronger memory effects at crossover
@@
Im[$\imaginary[\omega_{\text{FFC}}]>\Gamma_{\text{dec}}(\nu,\xi,\tau_{c}),$ (12)
@@
To quantify the potential suppression, we evaluate $\Gamma_{\text{dec}}$ using turbulence parameters from supernova simulations. Multi-dimensional studies show significant density fluctuations in the gain region, with relative amplitudes $\delta n_{e}/\bar{n}_{e}\sim 40\%$ yielding $\xi\sim 0.16$ [14]. The turbulence correlation scale in stratified regions like the proto-neutron star atmosphere is typically $L_{c}\sim 0.1-1$ km, giving $\tau_{c}\sim 0.1$ km in natural units [7]. For stratified turbulence, spectral exponents $\nu$ range from 1.0-1.0 (strong stratification) to 0.5-0.5 (moderate stratification) [6].
For the representative case $\nu=-1$, our exact solution gives $\Gamma_{\text{dec}}\sim\chi_{\nu}^{1/2}=2\sqrt{\xi\Delta_{m}\Gamma(2)}\approx 2.5\sqrt{\xi\Delta_{m}}$. The relevant oscillation frequency $\Delta_{m}$ in the gain region is dominated by the neutrino-neutrino potential $\mu_{\nu\nu}$, with typical values $\mu_{\nu\nu}\sim 10^{2}-10^{4}$ km-1. Taking $\Delta_{m}\sim 10^{3}$ km-1 and $\xi\sim 0.16$, we obtain:
$\Gamma_{\text{dec}}\sim 2.5\sqrt{0.16\times 10^{3}}\,\text{km}^{-1}\approx 32\,\text{km}^{-1}.$ (13)
This decoherence rate is comparable to typical FFC growth rates in the gain region ($\sim 10^{1}-10^{2}$ km-1) and can therefore significantly modify the instability threshold. Considering the range of possible $\mu_{\nu\nu}$ values ($10^{2}-10^{4}$ km-1), $\Gamma_{\text{dec}}$ spans 33003-300 km-1, which can suppress FFCs across much of the parameter space relevant to supernova conditions.
@@
For practical implementation, we propose that future supernova simulations extract turbulence parameters ($\nu$, $\xi$, $\tau_{c}$) along neutrino trajectories and apply the criterion Im[$\imaginary[\omega_{\text{FFC}}]>\Gamma_{\text{dec}}$ to determine FFC viability region-by-region. This provides a physics-based approach to assess whether and where FFCs might develop in realistic turbulent environments.
@@
The key physical discovery is that red turbulence induces accelerated decoherence: for the same integrated fluctuation strength $\xi$, flavor coherence persists longer than in white-noise environments but decays differently than in Kolmogorov-type turbulence. This decoherence behavior emerges from the algebraically growing temporal correlations $K(t)\sim t^{|\nu|}$ characteristic of stratified astrophysical plasmas.
Applying our results to the supernova fast flavor conversion problem, we find that the derived decoherence rate $\Gamma_{\text{dec}}(\nu,\xi,\tau_{c})$ can act as an effective damping mechanism in the linear stability analysis. For realistic parameters in supernova gain regions and proto-neutron star atmospheres ($\nu\sim-1$, $\xi\sim 0.1$, $\tau_{c}\sim 0.1$ km), this damping rate is $\sim 10$–100 km-1—comparable to typical FFC growth rates. Consequently, red turbulence can raise the FFC instability threshold by a factor of 3-5, potentially suppressing explosion-killing flavor equilibration in critical regions.